OpenStudy (anonymous):

What is the length of the transverse axis of the conic section shown below? (y+2)^2/16 - (x-3)^2/9 =1

OpenStudy (anonymous):

\[\frac{ (y+2)^2 }{ 16 } - \frac{ (x-3)^2 }{ 9 } =1\]

OpenStudy (anonymous):

would it be 2a

OpenStudy (jdoe0001):

for an ellipse, "a" is the BIGGER DENOMINATOR for a hyperbola, "a" is always in the POSITIVE fraction and yes, the traverse axis is 2a

OpenStudy (anonymous):

ohhh so 2a = 2(4) = meaning its 8 n_n

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

yay slowly i am getting the hang of this conics hahaha thank you @jdoe0001

OpenStudy (jdoe0001):

yw